A stone is dropped from a height of 25 meters. a) How long will it take for the stone to hit the ground? b) What velocity does it have when it hits the ground?

 



1
Expert's answer
2021-09-26T11:55:55-0400

Height, s=25m

Initial velocity, u=0

Acceleration due to gravity, g=10 m/s2

Final velocity, v=?

Time taken, t=?

(i) Using relation,

s=u t+1/2​gt2

25=0×t+1/2​×10×t2

25=0+5t2

t2=25/5​=5

t=square root of 5​=2.5 s

(ii)For a freely falling body:

v2=u2+2gh

=(0)2+2×(10)×(25)

So, v2=500

v= square root of 500​=25 m/s

 The speed of stone when it hits the ground will be 25m/s.